Evaluate the integral cos⁷3x dx with the limits from 0 to π/6.


1
Expert's answer
2021-05-24T14:38:01-0400

"\int _0^{\frac{\pi }{2}}\cos ^7\left(3x\right)\frac{1}{3}d3x"

"=\frac{1}{3}\cdot \int _0^{\frac{\pi }{2}}\cos ^7\left(3x\right)d3x"

"=\frac{1}{3}\cdot \int _0^{\frac{\pi }{2}}\left(1-\sin ^2\left(3x\right)\right)^3\cos \left(3x\right)d3x"

"=\frac{1}{3}\cdot \int _0^1\left(1-v^2\right)^3dv"

"=\frac{1}{3}\cdot \int _0^11-3v^2+3v^4-v^6dv"

"=\frac{1}{3}\left(\int _0^11dv-\int _0^13v^2dv+\int _0^13v^4dv-\int _0^1v^6dv\right)"

"=\frac{1}{3}\left(1-1+\frac{3}{5}-\frac{1}{7}\right)"

"=\frac{16}{105}"